OpenStudy (anonymous):

Find dy/dx, using implicit differentiation. lnx/y = 7 - x dx/dy =_____ Compare your answer with the result obtained by first solving for y as a function of x and then taking the derivative. y=____ dx/dy=_____ stuck on all 3 parts on this problem

OpenStudy (anonymous):

a little confused is this \[\ln(\frac{x}{y})\] or \[\frac{\ln(x)}{y}\]

OpenStudy (anonymous):

the 2nd one ln(x)/y

OpenStudy (anonymous):

then quotient rule for this one

OpenStudy (anonymous):

\[\frac{y\times \frac{1}{x}-\ln(x)y'}{y^2}=-1\]

OpenStudy (anonymous):

\[\frac{y}{x}-\ln(x)y'=-y^2\] \[-\ln(x)y'=-y^2-\frac{x}{y}\] \[y'=\frac{y^2}{\ln(x)}+\frac{x}{y\ln(x)}\] if my algebra is right

OpenStudy (anonymous):

kk so the first thing you gave me was the dy/dx= correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its not taking it for an answer :(

OpenStudy (rogue):

You switched up your x/y's Sat. It should be\[y'=\frac{y^2}{\ln(x)}+\frac{y}{x\ln(x)}\]

OpenStudy (anonymous):